# Propensity Score Alignment of Unpaired Multimodal Data

### v1

@@ Abstract
Multimodal representation learning techniques typically rely on paired samples to learn common representations, but paired samples are challenging to collect in fields such as biology where measurement devices often destroy the samples. This paper presents an approach to address the challenge of aligning unpaired samples across disparate modalities in multimodal representation learning. We draw an analogy between potential outcomes in causal inference and potential views in multimodal observations, which allows us to use Rubin’s framework to estimate a common space in which to match samples. Our approach assumes we collect samples that are experimentally perturbed by treatments, and uses this to estimate a propensity score from each modality, which encapsulates all shared information between a latent state and treatment and can be used to define a distance between samples. We experiment with two alignment techniques that leverage this distance—shared nearest neighbours (SNN) and optimal transport (OT) matching—and find that OT matching results in significant improvements over state-of-the-art alignment approaches in both a synthetic multi-modal setting and in real-world data from NeurIPS Multimodal Single-Cell Integration Challenge.
@@ 1 Introduction
Large-scale multimodal representation learning techniques such as CLIP (Radford et al., 2021) have lead to remarkable improvements in zero-shot classification performance and have enabled the recent success in conditional generative models. However, the effectiveness of multimodal methods hinges on the availability of paired samples—such as images and their associated captions—across data modalities. This reliance on paired samples is most obvious in the InfoNCE loss (Gutmann & Hyvärinen, 2010; van den Oord et al., 2018) used in CLIP (Radford et al., 2021) which explicitly learns representations to maximize the true matching between images and their captions.
@@
While paired image captioning data is abundant on the internet, paired multimodal data is often challenging to collect in scientific experiments. Take, for instance, experiments in biology, where unpaired data are the norm for technical reasons: RNA sequencing, protein expression assays, and the collection of microscopy images for cell painting experimental assays are all destructive processes. Because of this, we cannot collect multiple different measurements from the same cell, and can only explicitly group cells by their experimental condition.
If we could accurately match unpaired samples across modalities, we could use the aligned samples as proxies for paired samples and apply existing multimodal representation learning techniques. However, matching across modalities is challenging for two reasons. Firstly, observations from various modalities often exist in entirely different spaces. For instance, microscopy images of cells are in pixel space, while gene expression assays provide data in the form of mRNA abundance counts. Secondly, even if measurements were taken in the same space, determining an appropriate distance metric for matching remains a non-trivial task. For example, when matching images of distinct cells in different microscopy images, we would ideally want a metric that is more sensitive to features describing the underlying state of the cell than features related to biologically-irrelevant cell appearance, such as the orientation of the cell.
To understand the matching problem more abstractly, we can think of the multiple modalities as potential “views”, $x^{(1)}(z)\in\mathcal{X}^{(1)},x^{(2)}(z)\in\mathcal{X}^{(2)}$ of the same underlying latent state, $z\in\mathcal{Z}$, where we only get to observe one view for any individual unit $i$ (e.g. an individual cell in our motivating example). Observing, or accurately inferring $z$ would address the first issue by providing a common space $\mathcal{Z}$ in we could match samples. However, inferring the latent variable is hopelessly underspecified without making strong assumptions on the data generating processes of both modalities (Xi & Bloem-Reddy, 2023). Furthermore, for complex systems such as biological samples, $z$ may still be extremely high-dimensional, and as a result, the second challenge remains: what is the “right” metric is for measuring similarity between samples in this space?
@@
We address both challenges in this paper by appealing to classical ideas from Rubin (1974), in the case where we additionally observe a label for each unit $t\in\mathcal{T}$, e.g., indexing an experiment that it belongs to. By making the assumption that $t$ perturbs the observations via their shared latent state, we identify an observable link between modalities with the same underlying $z$. Our key observation is that the propensity score, defined as $p(t|z)$, satisfies three properties: 1) It provides a common space for matching, 2) it maximally coarsens $z$, retaining only relevant information and 3) under certain assumptions, it is estimable from observations of individual modalities alone (Proposition 3.1).
Figure 1: Our problem setting and proposed solution. We assume that two observed modalities (e.g., Microscopy images and RNA expression data) are generated in pairs according to some common latent variable, $z_{i}$, but we only observe the respective modalities separately. We further assume the latents are experimentally perturbed by a treatment, $t$, and we get observations from both modalities from a control state and at least one interventional state. Under our assumptions, training separate classifiers that predict the intervention label reveals the shared information $p(t|z_{i})$, which allows us to re-pair the observed disconnected modalities.
@@
In practice, the propensity score is straightforward to estimate: one simply trains two classifiers—one for each modality—to predict which treatment was applied to the respective modalities, and then match across modalities based on the similarity between predicted treatments within each treatment group. The proposed methodology is generic across modalities—it can be applied to match observations between any modality for which a classifier can be efficiently trained. However, we cannot naively use bipartite matching, as the same sample unit does not appear in both modalities. To address this, we use soft matching techniques to estimate the missing modality for each sample unit by combining multiple observations. We experiment with two different matching approaches from the recent literature: shared nearest neighbours (SNN) matching (Lance et al., 2022; Cao & Gao, 2022) and optimal transport (OT) matching Villani (2009); Schiebinger et al. (2019), and find that OT matching with distances defined on the propensity score lead to significant improvement on real world multi-modal matching tasks from the NeurIPS Multimodal Single-Cell Integration Challenge Lance et al. (2022). We found similarly large improvements on synthetic image-based benchmarks.
Finally, we show how our matching procedure can be used for cross-modality prediction tasks. We estimate a function that translates between modalities by integrating over the coupling matrix output by our alignment procedure. Our approach uses a two-sample stochastic gradient estimator to get an unbiased estimate of the gradient of the resulting loss. In our experiments, this approach gave very strong performance, even outperforming methods that had access to the ground truth matching.
2 Related work
Unpaired Data   Dealing with unpaired data is important in translation tasks, either for images (Liu et al., 2017; Zhu et al., 2017; Almahairi et al., 2018), or more recently for biological modalities (Amodio & Krishnaswamy, 2018; Yang et al., 2021). In particular, Yang et al. (2021) formalize this setting as being generated by a shared underlying latent variable, which has also been identified as a useful modelling assumption for identifiability analyses in multi-view or multi-modal representation learning (Gresele et al., 2020; Sturma et al., 2023). Our work also exploits this modelling assumption for theoretical justification, but, unlike previous works, our method does not depend on performing inference over the latent space, thus avoiding the identifiability issue. Specifically, since we do not require modality translation, we do not require an encoder-decoder structure (Yang et al., 2021), which embeds modality-specific reconstruction information in the latent space that can be detrimental for the simpler matching task.
Optimal Transport Matching   Optimal transport is commonly used to solve a similar problem in single-cell biology known as cell trajectory inference. In this case, $X^{(1)}$ to $X^{(2)}$ are random variables measured at different time points in a shared (metric) space, and OT matching can be achieved by minimizing the metric (Schiebinger et al., 2019; Tong et al., 2020). Recent work (Demetci et al., 2022) tackles our multi-modal alignment problem by using the Gromov-Wasserstein distance, which leverages local metric structure within pairs of points from each modality (Demetci et al., 2022). In addition to this “pure” OT approach, Gossi et al. (2023) use OT on contrastive learning representations, though this approach requires matched pairs for training, while Cao et al. (2022) use OT in the latent space of a multi-modal VAE.
@@ 2 Related work
Perturbations and Heterogeneity   Many methods treat observation-level heterogeneity as another dimension to globally integrate, even when cluster labels are observed (Butler et al., 2018; Korsunsky et al., 2019; Foster et al., 2022). This is sensible when clusters correspond to noise rather than the signal of interest. However, it is well known in causal representation learning that heterogeneity—particularly heterogeneity arising from different perturbations—has theoretical benefits (Khemakhem et al., 2020; Squires et al., 2023; Ahuja et al., 2023; Buchholz et al., 2023; von Kügelgen et al., 2023). There, the benefits (weakly) increase with the number of perturbations, which is also true of our setting (Proposition 3.2). In the context of aligning unpaired data, only Yang et al. (2021) leverage this heterogeneity in their method. Specifically, they require VAE representations to classify experimental labels in addition to reconstructing modalities, while our method only requires a representation to classify labels. Yang et al. (2021) treat the classifier objective as a regularizer, but our theory suggests that it is responsible for their matching performance. In our experiments, we found that requiring reconstruction led to worse matching performance with equivalent architectures.
3 Propensity Score Multi-modal Matching
@@
We consider multimodal settings where there exist two potential views, $X^{(e,t)}\in\mathcal{X}^{(e)}$ from two different modalities indexed by $e\in\{1,2\}$ and experiment $t$ that perturbs a shared latent state of these observations. The observations will typically be in very different spaces: for example, $\mathcal{X}^{(1,t)}$ may be the space of images of cells under a microscope, and $\mathcal{X}^{(2,t)}$ may be the space of gene expression data. This process defines a jointly distributed random variable $(X^{(1,t)},X^{(2,t)},e,t)$, from which we observe only a single modality, its index, and treatment assignment,$\{x_{i}^{(e_{i},t_{i})},e_{i},t_{i}\}_{i=1}^{n}$ (we will always denote random variables by upper-case letters, and samples by their corresponding lower-case letter). Our aim is to match or estimate the samples from the missing modality that would correspond to the realization of the missing random variable. Importantly, we match within treatment groups, and the treatment variable is only used to learn a common space in which to match.
@@
We assume each modality arises from a common latent random variable $Z$ as follows,
$t\sim P_{T},\quad Z^{(t)}\mid t\sim P_{Z}^{(t)},\quad U^{(e)}\sim P_{U}^{(e)},\quad U^{(e)}{\perp\!\!\!\perp}Z,$
@@
Under this model, if $Z$ were observable, an optimal matching can be constructed by simply matching the modalities with the most similar $Z$. However, $Z$ is latent, and inference on the model described by Section 3 is arguably more difficult than the matching problem itself due to theoretical difficulties such as identifiability and disentangling $Z$ from the modality-specific noise terms $U^{(e)}$.
Instead, we see that the interventions $t$ provide an observable link between the modalities, thereby revealing information about $Z^{(t)}$. Specifically, we use the propensity score with respect to $t$, which we define as
@@
as a proxy for the latent $Z$. Now, although we cannot compute this directly as $Z$ is latent, we make the observation that if $f^{(e)}$ is injective for $e=1,2$, then we can compute the compute the propensity score from each of the observed modalities, since it will be that $\pi(Z^{(t)})=\pi(X^{(e,t)})$, for both $e=1$ and $e=2$. Not only does the propensity score reveal shared information, classical causal inference (Rubin, 1974) states that it captures all information shared between the latent and treatment, and does so minimally, in terms of having minimum dimension and entropy. We collect these observations into the following proposition.
@@
In the model described by Section 3, further assume that $f^{(e)}$ are injective for $e=1,2$. Then, the propensity score in either modality is equal to the propensity score given by $Z^{(t)}$, i.e., $\pi(X^{(1,t)})=\pi(X^{(2,t)})=\pi(Z^{(t)})$ as random variables. This implies
$I(t,Z^{(t)}\mid\pi(Z^{(t)}))=I(t,Z^{(t)}\mid\pi(X^{(t)}))=0,$ (3)
for each $t$, where $I$ is the mutual information. Furthermore, any other function $b(Z^{(t)})$ satisfying $I(t,Z^{(t)}\mid b(Z^{(t)}))=0$ is such that $\pi(Z^{(t)})=f(b(Z^{(t)}))$.
The proof can be found in the Appendix. The above shows that computing the propensity score on either modality is equivalent to computing it on the unobserved shared latent, which captures all the shared information observable in $t$. The final statement implies that it is of minimal dimension and entropy, and thus it discards the modality-specific information that may be counterproductive to matching.
Number of Perturbations   Note that point-wise equality of the propensity score $\pi(z_{1})=\pi(z_{2})$ does not necessarily imply equality of the latents $z_{1}=z_{2}$, due to potential non-injectivity. In fact, suppose that $Z^{(t)}$ is supported in $\mathbb{R}^{d}$. The propensity score $\pi(z)$ maps into the $T$-dimensional simplex, where $T$ is the number of perturbations, which can be parameterized by coordinates in $\mathbb{R}^{T}$ (essentially, the logits of the classifier). Under mild conditions, the propensity score necessarily compresses information about $Z^{(t)}$ if the latent dimension exceeds the number of perturbations, echoing impossibility results from the causal representation learning literature (Squires et al., 2023).
@@
Suppose that $P_{Z}^{(t)}$ has a smooth density $p(z|t)$ for each $t=0,\dots,T$. Then, if $T<d$, the propensity score $\pi$, restricted to its strictly positive part, is non-injective.
@@
The proof can be found in the Appendix. Note the above only states an impossibility result when $T<d$. When $T\geq d$, it can be seen from the proof of Proposition 3.2 that the injectivity of the propensity score depends on the injectivity of the following expression in $z$:
$f(z)=\begin{bmatrix}\log(p(z|t=1))-\log(p(z|t=0))\\ \vdots\\ \log(p(z|t=T))-\log(p(z|t=0))\end{bmatrix},$ (4)

### v2

@@ Abstract
Multimodal representation learning techniques typically require paired samples to learn shared representations, but collecting paired samples can be challenging in fields like biology, where measurement devices often destroy the samples. This paper presents an approach to address the challenge of aligning unpaired samples across disparate modalities in multimodal representation learning. We draw an analogy between potential outcomes in causal inference and potential views in multimodal observations, allowing us to leverage Rubin’s framework to estimate a common space for matching samples. Our approach assumes experimentally perturbed samples by treatments, and uses this to estimate a propensity score from each modality. We show that the propensity score encapsulates all shared information between a latent state and treatment, and can be used to define a distance between samples. We experiment with two alignment techniques that leverage this distance—shared nearest neighbours (SNN) and optimal transport (OT) matching—and find that OT matching results in significant improvements over state-of-the-art alignment approaches in on synthetic multi-modal tasks, in real-world data from NeurIPS Multimodal Single-Cell Integration Challenge, and on a single cell microscopy to expression prediction task.
@@ 1 Introduction
Large-scale multimodal representation learning techniques such as CLIP (Radford et al., 2021) have lead to remarkable improvements in zero-shot classification performance and have enabled the recent success in conditional generative models. However, the effectiveness of multimodal methods hinges on the availability of paired samples—such as images and their associated captions—across data modalities. This reliance on paired samples is most obvious in the InfoNCE loss (Gutmann and Hyvärinen, 2010; van den Oord et al., 2018) used in CLIP (Radford et al., 2021) which explicitly learns representations to maximize the true matching between images and their captions.
Figure 1: Visualization of propensity score matching for two modalities (e.g., Microscopy images and RNA expression data). We first train classifiers to estimate the propensity score for samples from each modalities; the propensity score reveals the shared information $p(t|z_{i})$, which allows us to re-pair the observed disconnected modalities. The matching procedure is then performed within each perturbation class based on the similarity bewteen the propensity scores.
@@
While paired image captioning data is abundant on the internet, paired multimodal data is often challenging to collect in scientific experiments. For instance, unpaired data are the norm in biology for technical reasons: RNA sequencing, protein expression assays, and the collection of microscopy images for cell painting assays are all destructive processes. As such, we cannot collect multiple different measurements from the same cell, and can only explicitly group cells by their experimental condition. If we could accurately match unpaired samples across modalities, we could use the aligned samples as proxies for paired samples and apply existing multimodal learning techniques.
In this paper, we formalize this setting by viewing each modality as a potential measurement, $X^{(1)}(Z)\in\mathcal{X}^{(1)},X^{(2)}(Z)\in\mathcal{X}^{(2)}$, of the same underlying latent state $Z\in\mathcal{Z}$, where we are only able to make a single measurement for each sample unit (e.g. an individual cell). The task is to reconcile (match) unpaired observations $x^{(1)}$ and $x^{(2)}$ with the same (or maximally similar) $z$. Estimating the latent, $Z$, is hopelessly underspecified without making unverifiable assumptions on the system, and furthermore, $Z$ may still be sparse and high-dimensional, leading to inefficient matching. This motivates the need for approaches that use only the observable data.
We identify two major challenges for this problem. First, measurements are often made in very different spaces $\mathcal{X}^{(1)}$ and $\mathcal{X}^{(2)}$ (e.g., pixel space and gene expression counts), which make defining a notion of similarity across modalities challenging. Second, the measurement process inevitably introduces modality-specific variation that can be impossible to disentangle from the relevant information ($Z$). For example in cell imaging, we would not want the matching to depend on irrelevant appearance features such as the orientation of the cell or the lighting of the plate.
@@
In this paper, we address these challenges by appealing to classical ideas from causal inference (Rubin, 1974), in the case where we additionally observe some label $t$ for each unit, e.g., indexing an experiment. By making the assumption that $t$ perturbs the observations via their shared latent state, we identify an observable link between modalities with the same underlying $z$. Under conditions which we discuss in Section 2, the propensity score, defined as $p(t|Z)$, is a transformation of the latent $Z$ that satisfies three remarkable properties (Proposition 3.1): (1) it provides a common space for matching, (2) it is fully identifiable via classification on individual modalities, and (3) it maximally reduces the dimension of $Z$, retaining only the information revealed by the perturbations.
@@
The practical implementation of the methodology (as illustrated in Fig. 1) is then straightforward: we train two separate classifiers, one for each modality, to predict the treatment $t$ applied to $X^{(i)}$. We then match across modalities based on the similarity between the predicted probabilities (the propensity score) within each treatment group. This matching procedure is highly versatile and can be applied to match labeled observations between any modalities for which a classifier can be efficiently trained. However, since the same sample unit does not appear in both modalities, we cannot use naive bipartite matching. To address this, we use soft matching techniques to estimate the missing modality for each sample unit by allowing matching to multiple observations. We experiment with two recent matching approaches: shared nearest neighbours (SNN) matching (Lance et al., 2022; Cao and Gao, 2022) and optimal transport (OT) matching Villani (2009).
In our experiments, we find that OT matching with distances defined on the proposenity score leads to significant improvement on matching and a downstream cross-modality prediction task on both synthetic and real-world biological data. Notably, our prediction method, which leverages the soft matching to optimize an OT projected loss, outperforms supervised learning on the true pairs on CITE-seq data from the NeurIPS Multimodal Single-Cell Integration Challenge (Lance et al., 2022). Finally, we applied our method to match single-cell expression data (from a PeturbSeq assay (Dixit et al., 2016)) with single cell crops of image data (Fay et al., 2023). We find improved generalization in predicting the distribution of gene expression from the cell imaging data in with unseen perturbations.
1.1 Related Work
Unpaired and Multimodal Data
Learning from unpaired data has long been considered for image translation (Liu et al., 2017; Zhu et al., 2017; Almahairi et al., 2018), and more recently for biological modality translation (Amodio and Krishnaswamy, 2018; Yang et al., 2021). In particular, Yang et al. (2021) also takes the perspective of a shared latent variable for biological modalities. This setting has been studied more generally for multi-view representation learning (Gresele et al., 2020; Sturma et al., 2023) for its identifiability benefits.
@@ 1.1 Related Work
Perturbations and Heterogeneity   Many methods in biology treat observation-level heterogeneity as a nuisance dimension to globally integrate, even when cluster labels are observed (Butler et al., 2018; Korsunsky et al., 2019; Foster et al., 2022). This is sensible when clusters correspond to noise rather than the signal of interest. However, it is well known in causal representation learning that heterogeneity—particularly heterogeneity arising from perturbations—has theoretical benefits in constraining the solution set (Khemakhem et al., 2020; Squires et al., 2023; Ahuja et al., 2023; Buchholz et al., 2023; von Kügelgen et al., 2023). There, the benefits (weakly) increase with the number of perturbations, which is also true of our setting (Proposition 3.2). In the context of unpaired data, only Yang et al. (2021) explicitly leverage this heterogeneity in their method, while Ryu et al. (2024) treat it as a constraint in solving OT. Specifically, Yang et al. (2021) require their VAE representations to classify experimental labels in addition to reconstructing modalities, while our method is simpler, only requiring the classification objective. Notably, Yang et al. (2021) treat our objective as a regularizer, but our theory suggests that it is actually primarily responsible for the matching performance. Our experiment results coincide with the theoretical insights; requiring reconstruction, as in a VAE, led to worse matching performance with identical model architectures.
Optimal Transport Matching   OT is a common tool in single-cell biology. In cell trajectory inference, the unpaired samples are gene expression values measured at different time points in a shared (metric) space. OT matching minimizes this shared metric between time points (Schiebinger et al., 2019; Tong et al., 2020). Recent work (Demetci et al., 2022) extends this to our setting where each modality is observed in separate metric spaces by using the Gromov-Wasserstein distance, which computes the difference between the metric evaluated within pairs of points from each modality (Demetci et al., 2022). In concurrent work, this approach was recently extended to ensure matching within experimental labels (Ryu et al., 2024). In addition to these “pure” OT approaches, Gossi et al. (2023) use OT on contrastive learning representations, though this approach requires matched pairs for training, while Cao et al. (2022) use OT in the latent space of a multi-modal VAE.
2 Setting
@@
We consider the setting where there exist two potential views, $X^{(e)}\in\mathcal{X}^{(e)}$ from two different modalities indexed by $e\in\{1,2\}$, and experiment $t$ that perturbs a shared latent state of these observations. This defines a jointly distributed random variable $(X^{(1)},X^{(2)},e,t)$, from which we observe only a single modality, its index, and label, $\{x_{i}^{(e_{i})},e_{i},t_{i}\}_{i=1}^{n}$. [^1] We will denote random variables by upper-case letters, and samples by their corresponding lower-case letter. We aim to match or estimate the samples from the missing modality, which corresponds to the realization of the missing random variable. Since $t$ is observed, in practice we match observations within the same label class $t$.
@@
Formally, we assume each modality is generated by a common latent random variable $Z$ as follows:
$t\sim P_{T},\ Z^{(t)}\mid t\sim P_{Z}^{(t)},\ U^{(e)}\sim P_{U}^{(e)},\ U^{(e)}{\perp\!\!\!\perp}Z,\ U^{(e)}{\perp\!\!\!\perp}U^{(e^{\prime})},\ X^{(e)}\mid t=f^{(e)}(Z^{(t)},U^{(e)}),$ (1)
where $t$ indexes the experimental perturbations, and we take $t=0$ to represent a base environment. $U^{(e)}$ represents the modality-specific measurement noise that is unperturbed by $t$, and also independent across samples. The structural equations $f^{e}$ are deterministic after accounting for the randomness in $Z$ and $U$: it represents the measurement process that captures the latent state. For example, in a microscopy image, this would be the microscope and camera that maps a cell to pixels.
Comparison to Multimodal Generative Models
Our setting is technically that of a multimodal generative model with latent perturbations. However, by focusing on matching rather than generation, we are able to make significantly weaker and more meaningful assumptions while still ensuring the theoretical validity of our method. Without the effects of the perturbation, our Eq. 1 is essentially the same as (Yang et al., 2021, Equation 1) in an abstract sense. However, in order to fit the generative model, it is required to formulate explicit models over $f^{(e)}$ and $P_{Z}^{(t)}$, which requires specifying the function class (e.g., continuous) and the space of $Z$ (e.g., $\mathbb{R}^{d}$) as assumptions, even in universal approximation settings. In contrast, since we will not directly fit the model Eq. 1, we do not make any technical assumptions about the generative model. Instead, we will make the following assumptions on the underlying data generating process itself.
Key Assumptions   Our theory makes the following assumptions about the data generating process.
  • (A1)
@@
  • (A2)
    Injectivity of $f^{(e)}$: $f^{(e)}(z,u)=f^{(e)}(z^{\prime},u^{\prime})\implies(z,u)=(z^{\prime},u^{\prime})$. In words, each modality captures enough information to distinguish changes in the underlying state. [^2] Note that the injectivity is in the sense of $f$ as a function of both $u$ and $z$, which allows observations that have a shared $z$ but differ by their value in $u$, and the function remains injective. For example, rotated images with the exact same content can have a shared $z$, but remain injective due to the rotation being captured in $u$.
@@
3 Multimodal Propensity Scores
Under (1), if $Z$ were observable, an optimal matching can be constructed by simply matching the samples with the most similar $z_{i}$. However, the prerequisite of inverting the model and disentangling $Z$ is arguably more difficult than the matching problem itself. In particular, $Z$ is unidentifiable without strong assumptions on Eq. 1 (Xi and Bloem-Reddy, 2023), and even formulating the identifiability problem requires well-specification of the model as a prerequisite. We take an alternative approach that is robust to these problems, by using the perturbations $t$ as an observable link to reveal information about $Z$. Specifically, we show that the propensity score
@@ 3 Multimodal Propensity Scores
is identifiable as a proxy for the latent $Z$ under our assumptions of the data generating process. This is a consequence of the injectivity of $f^{(e)}$, since it will be that $\pi(Z)=\pi(X^{(e)})$, $e=1,2$, indicating that we can compute it from either modality. Not only does the propensity score reveal shared information, classical causal inference theory (Rubin, 1974) states that it captures all information about $Z$ that is contained in $t$, and does so minimally, in terms of having minimum dimension and entropy. Since $t$ contains the only observable information that is useful for matching, the propensity score is hence an optimal compression of the observed information. We collect these observations into the following proposition.
@@
In the model described by Eq. 1, further assume that $f^{(e)}$ are injective for $e=1,2$. Then, the propensity scores in either modality is equal to the propensity score given by $Z$, i.e., $\pi(X^{(1)})=\pi(X^{(2)})=\pi(Z)$ as random variables. This implies
$I(t,Z\mid\pi(Z))=I(t,Z\mid\pi(X^{(e)}))=0,$ (3)
for each $e=1,2$, where $I$ is the mutual information. Furthermore, any other function $b(Z)$ satisfying $I(t,Z\mid b(Z))=0$ is such that $\pi(Z)=f(b(Z))$.
The proof can be found in Appendix C. Practically, Proposition 3.1 shows that computing the propensity score on either modality is equivalent to computing it on the unobserved shared latent, which means that it is identifiable, and thus estimable, from the observations alone. Furthermore, the estimation does not require modified objectives or architectures for joint multimodal processing, instead they are simple and separate classification problems for each modality. Finally, $t$ does not affect $U^{(e)}$ by assumption, and thus the propensity score, being a representation of the information in $t$, discards the modality-specific information that may be counterproductive to matching. Therefore, even if $Z$ were observed, it may be sensible to match on its propensity score instead.
Number of Perturbations   Note that point-wise equality of the propensity score $\pi(z_{1})=\pi(z_{2})$ does not necessarily imply equality of the latents $z_{1}=z_{2}$, due to potential non-injectivity of $\pi$. For example, consider $t\in\{0,1\}$, then $\pi(z)$ is a compression to a single dimension $z\to p(t=1\mid z)$. Intuitively, collecting data from more perturbations improves the amount of information contained in the label $t$. If the latent space is $\mathbb{R}^{d}$, the propensity score necessarily compresses information about $Z^{(t)}$ if the latent dimension exceeds the number of perturbations, echoing impossibility results from the causal representation learning literature (Squires et al., 2023).
@@
Let $Z^{(t)}\in\mathbb{R}^{d}$. Suppose that $P_{Z}^{(t)}$ has a smooth density $p(z|t)$ for each $t=0,\dots,T$. Then, if $T<d$, the propensity score $\pi$, restricted to its strictly positive part, is non-injective.
@@
The proof can be found in Appendix C. Note the above only states an impossibility result when $T<d$. More generally, it can be seen from the proof of Proposition 3.2 that the injectivity of the propensity score depends on the injectivity of the following expression in $z$:
$g(z)=\begin{bmatrix}\log(p(z|t=1))-\log(p(z|t=0))\\ \vdots\\ \log(p(z|t=T))-\log(p(z|t=0))\end{bmatrix},$ (4)